Japan Company produces lamps that require 3 standard hours per unit at an hourly rate of $12.90 per hour. Production of 6,600 units required 20,390 hours at an hourly rate of $13.20 per hour. Enter favorable variances as negative numbers.

a. Determine the direct labor rate variance. Favorable or Unfavorable.

b. Determine the direct labor time variance. Favorable or Unfavorable.

c. Determine the cost variance.Favorable or Unfavorable.
Answer #1

Labor rate variance = (SR-AR) * AH

= (12.90 - 13.20) * 20,390

= 6117 UNFAVORABLE

.

Labor time variance = (SH-AH) * SR

= (6600*3 - 20390) * 12.90

= 7611 UNFAVORABLE

.

Labor cost variance = 6117 U + 7611 U

= 13,728 UNFAVORABLE